OpenStudy (anonymous):

Find the area of the regular polygon. Give the answer to the nearest tenth. decagon with a side of 4 cm A. 123.1 cm2 B. 129.4 cm2 C. 246.2 cm2 D. 139.8 cm2

ganeshie8 (ganeshie8):

Area of regular polygon of side length \(s\) : \(\large \dfrac{ns^2}{4} \cot (\frac{\pi}{n})\)

ganeshie8 (ganeshie8):

simply plugin \(s\), \(n\) values and evaluate

OpenStudy (anonymous):

n = number of sides s = length of side

OpenStudy (anonymous):

ohhh thanks for clarifying batman

OpenStudy (anonymous):

Np, so give this a shot, and get back to us :P

OpenStudy (anonymous):

Actually since we're using a decagon you can easily just use this: \[A = \frac{ 5 }{ 2 }a^2\sqrt{5+2\sqrt{5}}\] Where a is the side.

OpenStudy (anonymous):

40^2/4 =400

OpenStudy (anonymous):

decagon has 10 sides

OpenStudy (anonymous):

It's A !

OpenStudy (anonymous):

Correcto!